# MYTE: Morphology-Driven Byte Encoding for Better and Fairer Multilingual Language Modeling

### 15 Mar 2024 v1

@@ 2 Background: UTF-8 Bytes
Figure 2: UTF-8 codepage (inspired by the visualizations from: en.wikipedia.org/wiki/UTF-8). Each row contains 16 bytes with the same leading hexadecimal digit. Bytes in the range C2C2 - F4F4 are leading bytes. They mark the beginning of a multibyte code of the length shown in each cell. Bytes in the range 8080 - BFBF are continuation bytes, which follow a leading byte in multibyte codes. Bytes FEFE and FFFF are unused. Range C2C2 - F4F4 encodes Latin capital letters. In MYTE, these characters are decomposed to free space used to encode morphemes.
@@ 3.1 Morphological Analysis
0
Latin
Latin
42 4A 52
1
Common
Mixed, Common, Inherited, Unkown
43 4B 53
2
Non-Latin Alphabetic
Greek, Cyrillic, Armenian, Georgian
44 4C 54
3
Abjads
Hebrew, Arabic, Syriac, Thaana, Tifinagh
45 4D 55
4
Abugidas North
Devanagari, Gurmukhi, Gujarati, Oriya, Bengali, Sinhala, Tibetan
46 4E 56
5
Abugidas South
Telugu, Kannada, Tamil, Malayalam, Thai, Lao, Myanmar, Tai, Tagalog, Khmer
47 4F 57
6
CJK
Hangul, Han, Yi, Katakana, Hiragana, Bopomofo
48 50 58
7
Other
Remaining scripts
49 51 59
@@ 5.3 Results
Figure 6: Sentence prediction superisal expressed as Bit-per-English-Byte on multi-parallel Flores 200 corpus. The comparison shows that under MyT5 model, surprisal factors are more equitable than in the standard ByT5 model.
@@
As shown in Table 4, MyT5 and ByT5 perform comparably (and better than baselines) on MT and NER. While MyT5 outperforms ByT5 by 2 points on QA, the opposite is true for semantic parsing. We hypothesize that in this case, the morphological prior encoded in MYTE may confound semantic parsing fine-tuning, which requires a structured output significantly different from natural language.
@@ 6.1 Fair Representation across Languages
Some approaches aimed at balancing the segmentation or tokenization methods have been introduced. Limisiewicz et al. (2023) proposed merging vocabulary based on the tokenizer scoring function. Zheng et al. (2021) introduced a method of allocating vocabulary capacity uniformly across languages, while Chung et al. (2020) constructed multilingual vocabulary for clusters of languages and merged them. Liang et al. (2023) combined the elements of both approaches and showed the advantage of extending vocabulary to benefit multilingual transfer. These solutions offer the promise of obtaining a better allocation of vocabulary units. However, they do not solve the inequality of the underlying encoding, which may affect the construction process of vocabulary units. For instance, byte merges in the BPE algorithm always begin at individual bytes Sennrich et al. (2016); Zouhar et al. (2023).
@@ Appendix B Supplementary Results
Figure 8: Average byte sequence lengths of parallel sentences for languages unseen in the morphological analysis. Santali (sal) uses a script (Ol Chicki) distinct from all other languages included in morphological analysis.
@@ B.1 Results for Each Language
Figure 8 illustrates the sequence lengths and compressions obtained for languages unseen in the morphological analysis.
@@ B.2 LM Performance Across Scales
Figure 11 shows the difference of $BPEB$ between MyT5 and ByT5 the models in small and base scales. Furthermore, Table LABEL:tab:results_per_language contains average language modeling scores and inference times across all available scales for each language.
@@ C.2 Fine-Tuning
For few-shot fine-tuning, we choose the same hyperparameters and optimization strategy as in Ruder et al. (2023): $0.1$ dropout, $1e^{-3}$ learning rate with inverse square root decay. The batch size was chosen to facilitate training on v3-8 TPU, specifically 128 for NER; 64 for MT, QA, and semantic parsing. The number of fine-tuning steps corresponded to the sizes of the training datasets: QA 6500, NER 6000, semantic parsing 1000, and MT 10000.
For machine translation, we selected the following sample of language both for training and evaluation (only low-resource): Telugu, Tamil, Greek, Armenian, Russian, Kazakh, Amharic, Vietnamese, Japanese, French, Korean, German, Marathi, Polish.

### 11 Nov 2024 v2

@@ 2 Background: UTF-8 Bytes
Figure 2: UTF-8 codepage (inspired by the visualizations from: en.wikipedia.org/wiki/UTF-8). Each row contains 16 bytes with the same leading hexadecimal digit. Bytes in the range C2C2 - F4F4 are leading bytes. They mark the beginning of a multibyte code of the length shown in each cell. Bytes in the range 8080 - BFBF are continuation bytes, which follow a leading byte in multibyte codes. Bytes FEFE and FFFF are unused. Range 4141 - 5A5A encodes Latin capital letters. In MYTE, these characters are decomposed to free space used to encode morphemes.
@@ 3.1 Morphological Analysis
0 Latin Latin 4242 4A4A 5252
1 Common Mixed, Common, Inherited, Unkown 4343 4B4B 5353
2 Non-Latin Alphabetic Greek, Cyrillic, Armenian, Georgian $4$ 4C4C 5454
3 Abjads Hebrew, Arabic, Syriac, Thaana, Tifinagh 4545 4D4D $5$
4 Abugidas North Devanagari, Gurmukhi, Gujarati, Oriya, Bengali, Sinhala, Tibetan 4646 4E4E 5656
5 Abugidas South Telugu, Kannada, Tamil, Malayalam, Thai, Lao, Myanmar, Tai, Tagalog, Khmer 4747 4F4F 5757
6 CJK Hangul, Han, Yi, Katakana, Hiragana, Bopomofo 4848 5050 5858
7 Other Remaining scripts 4949 5151 5959
@@ 5.3 Results
Figure 6: Sentence prediction suprisal expressed as Bit-per-English-Byte on multi-parallel Flores 200 corpus. Each point corresponds to the BPEP value of one sentence. The comparison shows that under MyT5 model, performance is more equitable across languages than in the standard ByT5 model.
Figure 7: Bit-per-English-Byte for six languages unseen in morphological analysis: Santali, Sanskrit, Tunisian, Waray, Asturian, Esperanto. Santali (sal) uses an unseen script (Ol Chicki).
@@
In Figure 7, we observe that MyT5 outperforms ByT5 for languages unseen in morphological analysis, except for Sanatli, which also uses a distinct script.
@@
As shown in Table 4, MyT5 and ByT5 perform comparably (and better than baselines) on MT and NER. While MyT5 outperforms ByT5 by 2 points on QA, the opposite is true for semantic parsing. We hypothesize that in this case, the morphological prior encoded in MYTE may confound semantic parsing fine-tuning, which requires a structured output starkly dissimilar to natural language.
@@ 6.1 Fair Representation across Languages
Some approaches aimed at balancing the segmentation or tokenization methods have been introduced. Limisiewicz et al. (2023) proposed merging vocabulary based on the tokenizer scoring function. Zheng et al. (2021) introduced a method of allocating vocabulary capacity uniformly across languages, while Chung et al. (2020) constructed multilingual vocabulary for clusters of languages and merged them. Liang et al. (2023) combined the elements of both approaches and showed the advantage of extending vocabulary to benefit multilingual transfer. These solutions promise to obtain a better allocation of vocabulary units. However, they do not solve the inequality of the underlying encoding, which may affect the construction process of vocabulary units. For instance, byte merges in the BPE algorithm begin at individual bytes Sennrich et al. (2016); Zouhar et al. (2023). Therefore, the unequal granularity of UTF-8 representation impacts the vocabulary construction step in BPE, especially harming the low-resource non-Latin languages Kargaran et al. (2024). A possible solution is training BPE on top of MYTE encoded and balanced multilingual corpus.
@@ Appendix B Supplementary Results
Figure 9: Average byte sequence lengths of parallel sentences for languages unseen in the morphological analysis used in the construction of MYTE. Santali (sal) uses a script (Ol Chicki), a distinct script not seen in morphological analysis.
@@ B.1 Results for Each Language
Figure 9 illustrates the sequence lengths and compressions obtained for languages unseen in the morphological analysis. While Figure 7 shows the comparison of ByT5 and MyT5 for these languages.
@@ B.2 LM Performance Across Scales
Figure 12 shows the difference of $BPEB$ between MyT5 and ByT5 the models in small and base scales. Furthermore, Table LABEL:tab:results_per_language contains average language modeling scores and inference times across all available scales for each language. Both Figures show that MYTE offers improvement for languages not seen in morphological analysis but not for Sanatli, which uses a distinct script.
@@ C.2 Fine-Tuning
For few-shot fine-tuning, we choose the same hyperparameters and optimization strategy as in Ruder et al. (2023): $0.1$ dropout, $1e^{-3}$ learning rate with inverse square root decay. The batch size was chosen to facilitate training on v3-8 TPU, specifically 128 for NER; 64 for MT, QA, and semantic parsing. The number of fine-tuning steps corresponded to the sizes of the training datasets: QA 6500, NER 6000, semantic parsing 1000, and MT 10000. For machine translation, we selected the following sample of language both for training and evaluation: Telugu, Tamil, Greek, Armenian, Russian, Kazakh, Amharic, Vietnamese, Japanese, French, Korean, German, Marathi, and Polish.